OpenStudy (anonymous):

logbase4 8=x without a calculator

OpenStudy (unklerhaukus):

\[\log_48=\log_42^3=3\log_42\]

OpenStudy (unklerhaukus):

now use 4^{1/2}=2

OpenStudy (anonymous):

the answer is 3/2

OpenStudy (unklerhaukus):

good work

OpenStudy (anonymous):

i still dont understand how to do it

OpenStudy (unklerhaukus):

how did you arrive at 3/2 then?

OpenStudy (anonymous):

it's in the book. i lost you on your second step

OpenStudy (unklerhaukus):

\[3\log_42=3\log_44^{1/2}=\frac32\log_44\]

OpenStudy (anonymous):

thank you